A right triangle has one leg that is 2 ft. longer than the other leg. The hypotenuse is 2 ft. less than twice the shorter leg. Find the lengths of all sides of the triangle.
Answer by richwmiller(17219)

b=a+2,
a=b-2
c=2a-2,
a^2+b^2=c^2
a = 6, b = 8, c = 10